# Cobblestone: A Divide-and-Conquer Approach for Automating Formal Verification

### v3

@@ 3.4. RQ3: How Much Does Cobblestone’s Search Strategy Contribute to Its Performance?
Cobblestone’s search strategy is influenced by three key design decisions which we will examine in this RQ—(1) its use of whole-proof completions (2) its use of a recursive divide-and-conquer approach, and (3) its use of a sampling temperature of 1.0.
Cobblestone’s use of whole-proof completions sets it apart from many other proof synthesis systems. By contrast, most neural theorem provers synthesize proofs using tactic-by-tactic search, predicting the next tactic and using a tree search (typically depth-first or beam search) to search the tactic space (Yang19, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ; Sanchez20, ).
To measure the effectiveness of whole-proof completions, we implement another tool called TacticByTactic that, like Cobblestone, has use of the hammer, but uses GPT-4 to predict only the next tactic at each step of its search. While inefficient, TacticByTactic attempts the hammer tactic at each step of its search.
We run TacticByTactic, and TacticByTactic-NoHammer, a variant that does not use hammer, allowing up to 20 tactic predictions, with a maximum proof depth of 20, and 3 attempts to predict the next tactic at each step. As each prediction attempt results in one LLM sample, these tools use the same number of samples as Cobblestone.
@@
This datapoint is also interesting to observe because it is an ablation that runs fail-safe mode without recursively invoking Cobblestone. This ends up being similar to PALM’s (lu2024proofautomationlargelanguage, ) backtracking algorithm (which also does not make recursive calls).
As Cobblestone invokes itself recursively, using more LLM samples, the number of theorems it can prove increases even further. On CoqGym100, Cobblestone is able to prove 44 out of 48 (91.6%) theorems within 1 invocation, proving an additional 4 (9.4%) using recursive invocations. On Wigderson100, it is able to prove 31 out of 38 (81%) theorems within one invocation, proving 7 (19%) more over the course of its recursive calls. One possible explanation for this is test leakage in CoqGym100, which would make more samples of the original theorem “close enough” for Cobblestone’s to apply CoqHammer and error localization to produce a proof.
@@
The third distinguishing feature of Cobblestone’s search strategy is its use of temperature 1.0 when sampling whole proofs. Cobblestone benefits from sampling diverse proofs (recall Section 2.4). Setting the temperature to 1.0 contributes to this diversity. To examine the benefits of a higher temperature, we ran versions of Cobblestone and ChainOfThought with the temperature set to 0. Figure 11 shows the performance of these versions. In both Wigderson100 and CoqGym100, ChainOfThought-temp0 proves a subset of the theorems that ChainOfThought is able to prove. In Wigderson100, Cobblestone proves more theorems than Cobblestone-temp0, with a value added of 25%, and in CoqGym100, it proves the same number of theorems.
RA3: Cobblestone’s search strategy outperforms prior tools and an LLM-based tactic-by-tactic search. A single invocation of Cobblestone outperforms a combination of CoqHammer and ChainOfThought, and Cobblestone proves even more theorems when invoked recursively. A temperature of 1.0 can also improve performance.
@@ 3.5. RQ4: How Does External Information Affect Cobblestone’s Performance?
Cobblestone can use external information from another tool or a human. As a proxy, we created two oracles to provide such information (via the “API access points” in Figure 1. For each theorem, the perfect premises oracle (PerfPrems) knows the set of lemmas already proven in the project that the human-written proof for this theorem uses (e.g., with “rewrite” as described in Section 2.1). These lemmas are provided in addition to existing contextual information in the [OTHER PROVEN THEOREMS] section of the prompt in Section 2.4. Meanwhile, the perfect decomposition oracle (PerfDecomp) knows the set of subgoals the human-written proof proves. PerfDecomp provides a decomposing prefix (recall Section 2.5) from the human-written proof, helping fail-safe mode break the theorem into subgoals for recursive calls.
Using the PerfPrems oracle is equivalent to asking a human (or a tool) “What lemmas are relevant to proving this theorem?” and using the PerfDecomp oracle is equivalent to asking “How would you break down this theorem into smaller goals?”
We next evaluate Cobblestone with access to these oracles, calling the resulting variants Cobblestone-PerfPrems and Cobblestone-PerfDecomp. Cobblestone-PerfDecomp has access to both PerfPrems and PerfDecomp oracles. Figure 12 shows that Cobblestone-PerfPrems outperforms Cobblestone on both datasets, and Cobblestone-PerfDecomp sometimes performs even better. Interestingly, there is some complementarity to the variants, with each proving some theorems none of the others do. Together, the variants are able to prove 55% of CoqGym100 and 58% of Wigderson100. The additional information provided to Cobblestone by PerfDecomp does not always result in more theorems proven. Cobblestone-PerfPremises and Cobblestone are able to prove 6 theorems in CoqGym100 and 3 theorems in Wigderson100 that Cobblestone-PerfDecomp does not. This indicates that running Cobblestone several times, with different kinds of information in each run, may improve performance.
RA4: External information, such as useful already-proven lemmas or a decomposition of the theorem into subgoals can significantly increase Cobblestone’s proving power, suggesting a promising direction for future research.
@@ 3.6. RQ5: How do theorems Cobblestone proves and fails to prove differ?
To better understand where Cobblestone succeeds, we examine proofs that Cobblestone, Cobblestone-PerfPrems, and Cobblestone-PerfDecomp generate. To better understand when Cobblestone fails, we also examine the human-written proofs of some theorems where it fails. All proofs referenced in this section are available in our replication package(anonCobblestoneReplicationPackage, ).
@@
Successes: Cobblestone’s successful proofs range from 1 to 24 tactics in length, with the shortest consisting of a single invocation to CoqHammer. Successful proofs with length greater than 1 have an average length of 8.3 tactics, with bullet depth ranging from 1 (no bullets) to 4 (3 levels of bullets). These proofs can consist of anywhere from 1 to 9 distinct CoqHammer invocations or LLM samples. Of the proofs that decomposed their goal using bullets, prefixes ranged from 1 to 4 tactics long, with an average of 1.6. Of these prefixes, 41% decomposed using split, 36% used induction, 18% used destruct, and 14% used apply.
@@
The projects in our benchmarks broadly fall into two categories—mechanizations of mathematical proofs, and formalizations of software. CoqGym100 consists of 12 mathematical projects and 8 software projects, with 36 and 64 theorems in each category. Cobblestone proves 61.1% of CoqGym100’s mathematical theorems, and 40.6% of its software theorems. Additionally, coq-wigderson is a software project, and coq-bb5 and PnVRocqLib are mathematical projects. In total, our benchmarks contain 236 mathematical theorems and 164 software theorems, with Cobblestone proving 46.2% and 39.0% of these theorems respectively.
Together, Cobblestone-PerfPrems and Cobblestone-PerfDecomp generate 26 successful proofs for theorems unproven by Cobblestone (10 in CoqGym100 and 16 in Wigderson100). Interestingly, one of these proofs is of votesWithLog_update_elections_
data_timeout
@@
,333from RefinementSpecLemmas.v in verdi-raft whose human-written proof relies on break_match and tuple_inversion—custom tactics in Coq’s Ltac language. Cobblestone-PerfPrems generates an alternate proof without custom tactics. In the proof of Mcardinal_Scardinal,444from graph.v in coq-wigderson an LLM sample asserts a proposition that helps prove the theorem.
@@
Failures: Despite its improvements over the state-of-the-art, Cobblestone and its variants that use external information are unable to prove 45 theorems from Wigderson100 and 42 theorems from CoqGym100. We randomly sampled 14 such theorems (7 from each dataset) to examine manually.
@@
Many of these theorems’ human-written proofs are significantly longer than those that our tools generated proofs for. Cobblestone’s average proof is 10 tactics long, whereas 5 of the 14 unproven theorems have a human-written proof 20 or more tactics long. We also found that some theorems make it difficult for Cobblestone to take advantage of internal progress. For example, the ground-truth proof of requestVoteReply_term_sanity_client_
request555from RequestVoteReplyTermSanityProof.v in verdi-raft
@@
primarily consists of tactics like unfold and apply, each of which helps prove the theorem, with none breaking the goal into subgoals. It is difficult for Cobblestone to generate such proofs as it requires generating a working proof with a single LLM sample, preventing its divide and conquer algorithm from being able to recurse. Generalizing our approach to use partial progress that does not involve splitting a goal into subgoals may allow Cobblestone to prove more such theorems.
RA5: Cobblestone’s proofs are sometimes shorter than their human-written counterparts, and often successfully leverage its divide and conquer approach. Cobblestone’s failures indicate an opportunity to leverage new kinds of partial progress.
@@ 3.7. Threats to Validity
All LLM evaluations may suffer from test data leaking into the pretraining dataset, reducing generalizability to unseen data. We mitigate this risk by using coq-wigderson, coq-bb5, and PnVRocqLib, whose first GitHub commits are after GPT-4’s publicly stated pretraining cutoff date (recall Section 3.1). We evaluate on 400 theorems, in line with other published work (wanglego, ; wang2024proving, ; thakur2024an, ), to mitigate cost.
While our approach is general, our implementation is specific to Coq and uses GPT-4. We also evaluate with GPT-3.5-turbo and Claude 3 Opus to demonstrate generalizability.
Finally, our evaluation only began exploring how external information can aid Cobblestone’s proof synthesis and used oracles synthesized using human-written proofs. The intent of these oracles is to proxy how a human might interact with Cobblestone, but are too perfect to make any predictions on how an interactive, semi-automated approach may perform. More research is necessary to study the effects of incomplete and noisy data, as well as user studies to show the potential impact of real, human-provided information.
@@ 4. Related Work
Recent work in automating theorem proving in proof assistants has mostly explored three overarching approaches: hammers, machine learning techniques, and combinations of the two. Hammers, such as CoqHammer (Czajka18, ) and Sledgehammer (Paulson23, ) for Isabelle, call out to SMT solvers (schulz2013system, ; deMoura08, ) to construct low-level proofs, though they cannot use induction, and so are limited in what they can prove on their own. Our evaluation has shown that our approach can often prove theorems CoqHammer fails to prove on its own.
@@
Machine learning techniques called neural theorem provers typically use a predictive model learned from a corpus of existing proofs to predict the next steps, and use these predictions to guide a search through the space of potential proofs (li2024survey, ). They have been built using RNNs (Huang18, ; Sanchez20, ), LSTMs (Yang19, ; First20oopsla, ; First22icse, ), GNNs (bansal2019learning, ; Blaauwbroek24ICML, ), and transformer-based LLMs (Jiang21, ; Yang23, ). LLM-based methods prompt the pretrained model zero-shot or with few-shot examples (jiang2023draft, ; zhang2023getting, ), fine-tune the model on proof data (First23fse, ; huang2024mustard, ), or use it as an agent (thakur2023languageagent, ; wanglego, ).
@@
Hammers and machine learning techniques are complementary (Jiang2022Thor, ; First22icse, ). Thor (Jiang2022Thor, ) fine-tunes an LLM to learn when to apply Sledgehammer to a subgoal or when to predict a tactic. By contrast, Cobblestone samples whole proofs rather than tactics and makes hammer calls without prediction. Focusing on formalizing math, DraftSketchProve (jiang2023draft, ) uses an LLM to translate informal proofs into a formal proof sketch with holes, filling in the holes with calls to Sledgehammer, with others extending this framework (zheng2023lyra, ; wanglego, ; zhao2023decomposing, ). LEGO (wanglego, ) decomposes theorems into helper lemmas using an LLM, providing no guarantee that proving them will prove the original theorem; by contrast, Cobblestone uses the theorem prover to decompose goals with such a guarantee.
@@
LeanDojo (Yang23, ) and Magnushammer (Mikula24, ) train models to select relevant premises (lemmas/definitions) at each proof step. In addition to selecting relevant premises, Rango (thompson2025rango, ) also selects similar proof scripts. Cobblestone uses the preceding lemmas in the file, but could benefit from a premise selection model or a proof-script selection model, in the same way it benefits from external information in the form of an oracle’s perfect premises.
@@
Proof engineers often repair previously working proofs (Ringer20, ). Symbolic tools can automate some repair (Ringer21, ; masci2022proof, ). Baldur uses error messages to fine-tune an LLM to repair Isabelle proofs, in the context of proof synthesis (First23fse, ). Unlike Baldur, which only attempts one repair, Cobblestone is an iterative, divide-and-conquer approach. PALM (lu2024proofautomationlargelanguage, ) also samples from an LLM and uses repair mechanisms. However, as detailed in Section 3.4, our approach is recursive in nature.
@@
POETRY fine-tunes an LLM and estimates proof points on which to recurse (wang2024proving, ). By contrast, Cobblestone requires no tuning training data and uses fail-safe mode to precisely identify subgoals that recursion. Adapting Cobblestone to use more powerful models requires only minor changes to its API calls, whereas adapting POETRY is only possible with access to the model’s weights, and would require significant resources for fine-tuning.
@@
Formally specifying system properties is complementary to proving them. Prior work has tackled formalizing natural language specifications (Endres24, ; Goffi16, ; Motwani19icse, ; Zhai20, ). Once specified, some properties, like fairness (Galhotra17fse, ), can be verified probabilistically (Thomas19science, ; Hoag23icse-demo, ; Metevier19neurips, ; Giguere22iclr, ; Albarghouthi17, ).
@@
Prompt engineering is an important aspect of effective LLM use (openai2023gpt4, ; chowdhery2023palm, ; touvron2023llama, ). Prompting frameworks, such as chain-of-thought (wei2022chain, ) and others (yao2023tree, ; chen2024boosting, ; besta2023graph, ), can elicit reasoning in LLMs (wei2022chain, ; wang2022self, ; zhou2022least, ; fu2022complexity, ; chu2023survey, ). LLMs can be used for quantitative reasoning tasks (Lewkowycz2022Minerva, ; azerbayev2023llemma, ) and programming tasks (rozière2023code, ; chen2021evaluating, ). These advances are complementary to our efforts to generate proofs automatically.

### v4

@@ 3.4. RQ3: How Much Does Cobblestone’s Search Strategy Contribute to Its Performance?
Cobblestone’s search strategy is influenced by three key design decisions which we examine in this research question—(1) its use of whole-proof completions (2) its use of a recursive divide-and-conquer approach, and (3) its use of a sampling temperature of 1.0.
Cobblestone’s use of whole-proof completions sets it apart from many other proof synthesis systems. By contrast, most neural theorem provers synthesize proofs using tactic-by-tactic search, predicting the next tactic and using a tree search  (Yang and Deng, 2019; First et al., 2020; First and Brun, 2022; Sanchez-Stern et al., 2023, 2020).
@@
To measure the effectiveness of whole-proof completions, we implement another tool called TacticByTactic that, like Cobblestone, uses hammer (one for each search step), but only uses GPT-4 to predict the next tactic at each search step. We run TacticByTactic, and TacticByTactic-NoHammer, a variant that does not use hammer, allowing up to 20 tactic predictions, a maximum proof depth of 20, and 3 attempts to predict the next tactic at each step. As each prediction attempt results in one LLM sample, these tools use the same number of samples as Cobblestone.
@@
This datapoint is also interesting to observe because it is an ablation that runs fail-safe mode without recursively invoking Cobblestone. This ends up being similar to PALM’s (Lu et al., 2024) backtracking algorithm (which also does not make recursive calls).
Invoking Cobblestone recursively further increases the number of theorems it can prove. On CoqGym100, Cobblestone is able to prove 44 out of 48 (91.6%) theorems within 1 invocation, proving an additional 4 (9.4%) by recursing. On Wigderson100, it is able to prove 31 out of 38 (81%) theorems within one invocation, proving 7 (19%) more by recursing. One possible explanation for this is test leakage in CoqGym100, which would make more samples of the original theorem “close enough” for Cobblestone’s to apply CoqHammer and error localization to produce a proof.
@@
The third distinguishing feature of Cobblestone’s search strategy is its use of temperature 1.0 when sampling whole proofs. Cobblestone benefits from sampling diverse proofs (recall Section 2.4). A temperature of 1.0 contributes to this diversity. To examine the benefits of a higher temperature, we ran Cobblestone and ChainOfThought with the temperature set to 0. As shown in Figure 12, in both Wigderson100 and CoqGym100, ChainOfThought-temp0 proves a subset of the theorems that ChainOfThought is able to prove. In Wigderson100, Cobblestone proves more theorems than Cobblestone-temp0, with a value added of 25%, and in CoqGym100, it proves the same number of theorems.
[…]
@@ 3.5. RQ4: How Does External Information Affect Cobblestone’s Performance?
Cobblestone can use external information from another tool or a human. We created two oracles to provide such information (via the “API access points” in Figure 1). For each theorem, the perfect premises oracle (PerfPrems) knows which proven lemmas the human-written proof uses (e.g., with “rewrite” as described in Section 2.1). These are provided in addition to existing context in the [OTHER PROVEN THEOREMS] section of the prompt in Section 2.4. The perfect decomposition oracle (PerfDecomp) provides a decomposing prefix (recall Section 2.5) from the human-written proof, breaking the theorem into subgoals for recursive calls. Using the PerfPrems oracle is equivalent to asking “What lemmas are relevant to proving this theorem?” and using the PerfDecomp oracle is equivalent to asking “How would you break down this theorem into smaller goals?”
We evaluate Cobblestone with access to these oracles, calling the resulting variants Cobblestone-PerfPrems and Cobblestone-PerfDecomp. Cobblestone-PerfDecomp has access to both PerfPrems and PerfDecomp oracles. Figure 13 shows that Cobblestone-PerfPrems outperforms Cobblestone on both datasets, and Cobblestone-PerfDecomp sometimes does even better. Interestingly, there is some complementarity to the variants, with each proving some theorems none of the others do. Together, the variants prove 55% of CoqGym100 and 58% of Wigderson100. The additional information provided to Cobblestone by PerfDecomp does not always result in more theorems proven. Cobblestone-PerfPremises and Cobblestone are able to prove 6 theorems in CoqGym100 and 3 theorems in Wigderson100 that Cobblestone-PerfDecomp does not. This indicates that running Cobblestone several times, with different kinds of information in each run, may improve performance.
@@
[…]
@@ 3.6. RQ5: How do theorems Cobblestone proves and fails to prove differ?
To better understand where Cobblestone succeeds, we examine proofs that Cobblestone and its oracle-augmented variants generate. To better understand when it fails, we also examine human-written proofs of some theorems it cannot prove. All proofs referenced are available in our replication package(Kasibatla et al., 2025).
@@
Successes: Cobblestone’s successful proofs range from 1 to 24 tactics long, with the shortest consisting of a single invocation to CoqHammer. Successful proofs with length greater than 1 have an average length of 8.3 tactics, with bullet depth ranging from 1 (no bullets) to 4 (3 levels of bullets). These proofs can consist of 1 to 9 distinct CoqHammer invocations or LLM samples. Of the proofs that decomposed their goal, prefixes ranged from 1 to 4 tactics long, with an average of 1.6. Of these prefixes, 41% decomposed using split, 36% used induction, 18% used destruct, and 14% used apply.
@@
The projects in our benchmarks broadly fall into 2 categories—mechanizations of mathematical proofs and formalizations of software. CoqGym100 consists of 12 mathematical projects and 8 software projects with 36 and 64 theorems respectively. Cobblestone proves 61.1% of CoqGym100’s mathematical theorems and 40.6% of its software theorems. coq-wigderson is a software project, and coq-bb5 and PnVRocqLib are mathematical projects. Including these, our benchmarks contain 236 mathematical and 164 software theorems, with Cobblestone proving 46.2% and 39.0% respectively.
Together, Cobblestone-PerfPrems and Cobblestone-
@@
PerfDecomp generate 26 successful proofs for theorems unproven by Cobblestone (10 in CoqGym100 and 16 in Wigderson100). Interestingly, in one of these proofs555votesWithLog_update_elections_data_timeout from RefinementSpecLemmas.v in verdi-raft, the human-written proof relies on custom tactics in Coq’s Ltac language, while Cobblestone-PerfPrems generates a proof without custom tactics. In another [^6] Mcardinal_Scardinal from graph.v in coq-wigderson, Cobblestone asserts a proposition that helps prove the theorem.
@@
Failures: Despite its improvements over the state-of-the-art, Cobblestone and its variants cannot prove 45 theorems from Wigderson100 and 42 from CoqGym100. We randomly sampled 14 such theorems (7 from each dataset) to examine manually.
@@
Many of these theorems’ human-written proofs are much longer than Cobblestone’s proofs. Cobblestone’s average proof is 10 tactics long, whereas 5 of the 14 unproven theorems have a human-written proof of 20+ tactics. We also found that some theorems make it difficult for Cobblestone to make partial progress. For example, the ground-truth proof of request_VoteReply_term_
sanity_client_request777from RequestVoteReplyTermSanityProof.v in verdi-raft
@@
uses tactics like unfold and apply, each of which helps prove the theorem, without decomposing the goal. It is difficult for Cobblestone to generate such proofs as it requires generating a working proof without recursing. Generalizing our approach to use partial progress that does not involve splitting a goal into subgoals may allow Cobblestone to prove such theorems.
[…]
@@ 3.7. Threats to Validity
All LLM evaluations may suffer from test data leaking into the pretraining dataset, reducing generalizability to unseen data. We mitigate this risk by using coq-wigderson, coq-bb5, and PnVRocqLib, whose first GitHub commits are after GPT-4’s publicly stated pretraining cutoff date (recall Section 3.1). We evaluate on 400 theorems, in line with other published work (Wang et al., 2024b, a; Thakur et al., 2024a), to mitigate cost.
While our approach is general, our implementation is specific to Coq and uses GPT-4. We also evaluate with GPT-3.5-turbo and Claude 3 Opus to show generalizability. When evaluating against other LLM-based approaches such as PALM (Lu et al., 2024) and Rango (Thompson et al., 2025), we give each tool similar resources e.g. by running PALM 15x to control for tokens. However, Cobblestone and Rango’s token consumption is not straightforward to compare, as Rango uses a model with fewer parameters and a different tokenization algorithm.
Finally, our evaluation only began exploring how external information can aid Cobblestone’s proof synthesis using oracles with access to human-written proofs. The intent of these oracles is to proxy how a human might interact with Cobblestone. But, their performance does not indicate how an interactive, semi-automated approach may perform. More research is necessary to study the effects of incomplete and noisy data, as well as user studies to show the potential impact of real, human-provided information.
@@ 4. Related Work
Recent work automating theorem proving in proof assistants has mostly explored three overarching approaches: hammers, machine learning techniques, and combinations of the two. Hammers such as CoqHammer (Czajka and Kaliszyk, 2018) and Sledgehammer (Paulson and Nipkow, 2023) call SMT solvers (Schulz, 2013; de Moura and Bjørner, 2008) to construct low-level proofs. They cannot use induction, and thus are limited in what they can prove. Our evaluation shows that Cobblestone proves many theorems CoqHammer cannot.
@@
Machine learning techniques (i.e. neural theorem provers) use a predictive model learned from existing proofs to predict next steps, and use these predictions to guide a search through the space of proofs (Li et al., 2024). They have been built using RNNs (Huang et al., 2018; Sanchez-Stern et al., 2020), LSTMs (Yang and Deng, 2019; First et al., 2020; First and Brun, 2022), GNNs (Bansal et al., 2019; Blaauwbroek et al., 2024), and LLMs (Jiang et al., 2021; Yang et al., 2023). LLM-based methods prompt a pretrained model zero-shot or with few-shot examples (Jiang et al., 2023; Zhang et al., 2023), fine-tune the model (First et al., 2023; Huang et al., 2024), or use it as an agent (Thakur et al., 2024b; Wang et al., 2024b).
@@
Hammers and machine learning techniques are complementary (Jiang et al., 2022; First and Brun, 2022). Thor (Jiang et al., 2022) fine-tunes an LLM to learn when to apply Sledgehammer vs. predict a tactic. By contrast, Cobblestone samples whole proofs rather than tactics and calls hammer without fine-tuning. Focusing on formalizing math, DraftSketchProve (Jiang et al., 2023) uses an LLM to translate informal proofs into formal proof sketches with holes, filling the holes with Sledgehammer. Others extend this framework (Zheng et al., 2024; Wang et al., 2024b; Zhao et al., 2024). LEGO (Wang et al., 2024b) decomposes theorems into helper lemmas using an LLM, but provides no guarantee that proving them will prove the original theorem; by contrast, Cobblestone uses the theorem prover to decompose goals with such a guarantee.
@@
LeanDojo (Yang et al., 2023) and Magnushammer (Mikuła et al., 2024) train models to select relevant premises (lemmas/definitions) at each proof step. In addition to premises, Rango (Thompson et al., 2025) also selects similar proof scripts. Cobblestone uses the preceding lemmas in the file, but could benefit from a premise selection or a proof-script selection model in the same way it benefits from perfect premises provided by an oracle.
@@
Proof engineers often repair previously working proofs (Ringer et al., 2020). Symbolic tools can automate some repair (Ringer, 2021; Masci and Dutle, 2022). Baldur fine-tunes an LLM to repair Isabelle proofs using error messages (First et al., 2023). Unlike Baldur, which only attempts one repair, Cobblestone is an iterative, divide-and-conquer approach. PALM (Lu et al., 2024) samples from an LLM and uses repair mechanisms. By contrast, as detailed in Section 3.4, our approach is recursive in nature.
@@
POETRY fine-tunes an LLM and estimates proof points on which to recurse (Wang et al., 2024a). By contrast, Cobblestone requires no tuning training data and uses fail-safe mode to precisely identify subgoals that recursion. Adapting Cobblestone to use more powerful models requires only minor changes to its API calls, whereas adapting POETRY is only possible with access to the model’s weights, and would require significant resources for fine-tuning.
@@
Formally specifying system properties is complementary to proving them. Prior work has tackled formalizing natural language specifications (Endres et al., 2024; Goffi et al., 2016; Motwani and Brun, 2019; Zhai et al., 2020). Once specified, some properties, like fairness (Galhotra et al., 2017), can be verified probabilistically (Thomas et al., 2019; Hoag et al., 2023; Metevier et al., 2019; Giguere et al., 2022; Albarghouthi et al., 2017).
@@
Prompt engineering is key to effective LLM use (OpenAI, 2023; Chowdhery et al., 2023; et. al., 2023). Frameworks, such as chain-of-thought (Wei et al., 2022b) and others (Yao et al., 2023; Chen et al., 2024; Besta et al., 2024), can elicit reasoning in LLMs (Wei et al., 2022b; Wang et al., 2023; Zhou et al., 2023; Fu et al., 2023; Chu et al., 2024). LLMs can be used for quantitative reasoning tasks (Lewkowycz et al., 2022; Azerbayev et al., 2024) and programming tasks (Rozière et al., 2023; Chen et al., 2021). These advances complement our efforts to generate proofs automatically.